OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

rearrange -3y = -12-4x then y = (-12-4x)/-3 so y = 4 + (3/4)x so by comparing with y=mx+C then m is the gradient which is 3/4 gradient of the parallel will have same gradient. Now substitute y and x from the point given and m (3/4) into the line equation y=mx+C find C by rearranging then replace m and C in the line equation and you have your answer. This is the solution that I spoke about x=-3 and y=4 and m = 3/4 so replacing in equation 4=-3x3/4 +C so C = 4+3x3/4 =4+9/4 = (16+9)/4=25/4 now substitute in equation y= (3/4)x + 25/4 if you multiply by 4 equation looks better 4y = 3x+25